# xLSTM Scaling Laws: Competitive Performance with Linear Time-Complexity

### v1

@@ 7 Conclusion
Our study provides a systematic comparison of scaling behaviors between xLSTM and Transformer architectures. We show that xLSTMs are Pareto-dominant in training loss versus compute, maintain consistent power-law exponents in the overtraining regime, and scale more efficiently with context length due to their linear complexity. While our results suggest a universal relationship between performance and model size that applies to both compute-optimal Transformers and xLSTM models, we find that compute-optimal xLSTM models are larger than their Transformer counterpart and that the compute-optimal model size of xLSTMs is robust to variations in context length. During inference, xLSTM models achieve lower time to first tokens and generation step times than Transformer models of the same size. These results are well explained by our runtime model, which is grounded in theoretical FLOP and memory operation calculations and shows close agreement with the empirical data. Throughout all experiments, we find that the advantages of xLSTM grow with context length, both for training and inference characteristics, positioning xLSTM as a promising and scalable architecture for future language models.
@@ A.1 Details on the Experimental Setup
In Appendix E we provide a list of model architecture configurations for all Transformer and xLSTM models used in our scaling law study in Token/Param (App. E.1) and IsoFLOP (App. E.2) training setups.
@@ A.4 Additional Results: IsoFLOP Approach
In the main paper (Sec. 3.4, Fig 4), we presented results for the compute-optimal model size. In Fig. 9 we present results w.r.t. the number of training tokens. We observe that compute-optimal xLSTMs and Transformers are trained on a similar number of tokens.
@@ A.5 Additional Results: IsoFLOP Approach for Different Context Lengths
Appendix B Code and Reproducability
We release the code to reproduce our figures and the datasets of training runs in our public code repository (https://github.com/NX-AI/xlstm_scaling_laws) to facilitate future research in this direction. The datasets of training runs have been obtained using the publicly available xLSTM 7B training repository (https://github.com/NX-AI/xlstm-jax) using the model configurations stated in Appendix E. Inference results have been obtained using the publicly available benchmarking pipeline for efficient xLSTM kernels (https://github.com/NX-AI/mlstm_kernels), more specifically, the model benchmarks, not those for individual kernels.
Appendix C Accounting: Parameters, Cache Sizes, FLOPs, Memory Operations
In this section, we count number of parameters (App. C.1), memory state or KV cache size C.2, FLOPs (App. C.3), and memory operations (App. C.4) for mLSTM models based on the architecture of xLSTM 7B (Beck et al., 2025b) and Transformer models with Self-Attention based on the Llama 3 architecture (Grattafiori et al., 2024).
@@ Appendix C Accounting: Parameters, Cache Sizes, FLOPs, Memory Operations
C.1 Parameter Counts
We count the number of parameters of mLSTM models (C.1.1) and Transformer models (C.1.2). We include embedding and normalization layer parameters in our parameter counts.
C.1.1 mLSTM Params
@@ C.1.2 Transformer Params
C.2 Memory State and KV-Cache Size
@@
C.3 FLOP Counts
In this section, we count the FLOPs for the mLSTM and the Transformer model architecture. For each model architecture we count the sequence length dependent FLOPs for the sequence mix layer first, i.e. the mLSTM cell (C.3.1) and the Self-Attention layer (C.3.3), and then combine them with the FLOPs of the other layers in the model architecture to obtain the total FLOPs for the mLSTM (C.3.2) and the Transformer model (C.3.4).
We do not drop subleading terms and set also count FLOPs for all operations equally, i.e. $F_{\text{OP}}=1$. We also count the FLOPs for the normalization layers with $F_{\text{norm}}=3$ (we assume the factor of 3 because we have mean, variance and division operations). The skip connection FLOPs are counted with $F_{\text{skip}}=1$, or if neglected with $F_{\text{skip}}=0$.
C.3.1 mLSTM Cell FLOPs
@@ C.3 FLOP Counts
C.3.2 mLSTM Model FLOPs
The number of FLOPs for the backbone is identical for training, prefill and generation as the operations (embeddings, linear layers and layernorms) do not depend on the sequence length. Therefore, we count the FLOPs per token for the mLSTM backbone. To obtain the total FLOPs for the specific setting we have to use the respective expression for the mLSTM cell FLOPs from Appendix C.3.1.
@@
C.3.3 Self-Attention FLOPs
@@
C.3.4 Transformer Model FLOPs
Similar to the mLSTM backbone in Appendix C.3.2, the number of FLOPs for the Transformer backbone is identical for training, prefill and generation as the operations (embeddings, linear layers and layernorms) do not depend on the sequence length. Therefore, we count the FLOPs per token for the Transformer backbone. To obtain the total FLOPs for the specific setting we have to use the respective expression for the Self-Attention layer FLOPs from Appendix C.3.3.
@@ C.3.4 Transformer Model FLOPs
C.4 Memory Operation Counts
In this section, we count the memory operations for the mLSTM and the Transformer model architecture. We follow the same outline as for the FLOP counts in Appendix C.3 and first count the memory operations for the mLSTM cell (C.4.1) and the Self-Attention layer (C.4.3), and then combine them with the memory operations of the other layers in the model backbone to obtain the total memory operations for the mLSTM (C.4.2) and the Transformer model (C.4.4). We model weight MemOps as a one-time streaming cost (perfect on-chip reuse), i.e., independent of the number of token in the batch $BT$. This is reasonable with persistent/fused kernels and per-rank weight matrices that fit in on-chip cache. Depending on the exact experimental configuration, this assumption might not hold as we observe when modeling the step time through MemOps in Section D.3.
@@ C.4 Memory Operation Counts
C.4.1 mLSTM Cell MemOps
Similar to the FLOP counts in Appendix C.3.1, we count the memory operations for the mLSTM cell for both the chunkwise-parallel and the recurrent formulation.
@@
C.4.2 mLSTM Model MemOps
@@
C.4.3 Self-Attention MemOps
Similar to the FLOP counts in Appendix C.3.3, we count the memory operations for a single Self-Attention head during training or prefill and generation.
@@ C.4.3 Self-Attention MemOps
Similar to the FLOP counts in Appendix C.3.3, also the memory operation counts for the Self-Attention layer during generation depend on the number of previous tokens $T=T_{\text{p}}+t_{\text{g}}$, where $T_{\text{p}}$ is the number of prefill tokens and $t_{\text{g}}$ is the number of generated tokens so far.
@@
C.4.4 Transformer Model MemOps
Similar to the mLSTM backbone in Appendix C.4.2, the number of memory operations for the Transformer backbone (excluding the Self-Attention layer) consist of the input and output activations as well as the parameters. The memory operations for input and output activations depend on the number of tokens $BT$ in the batch, whereas the parameters are independent of the number of tokens.
@@ C.4.4 Transformer Model MemOps
Appendix D Modeling Inference Characteristics
In this section, we create a model of the theoretical runtimes of operations in the xLSTM and Transformer model architectures to model their inference characteristics (TTFT and step time). This theoretical model is based on the FLOP and the memory operation counts in Appendix C.
@@ Appendix D Modeling Inference Characteristics
D.1 Background: Theoretical Runtime
@@
For accelerator speed $\alpha_{\text{ acc}{{}}}$, accelerator memory bandwidth $\beta_{\text{ acc}{{}}}$, and accelerator communication bandwidth $\gamma_{\text{ Bytes}{{}}}$, we use the hardware specifications of NVIDIA V100111https://www.nvidia.com/en-au/data-center/v100/, A100222https://www.nvidia.com/en-us/data-center/a100/, H100333https://www.nvidia.com/en-au/data-center/h100/ and B200444https://resources.nvidia.com/en-us-blackwell-architecture/datasheet GPUs, which we summarize in Tab. 18.
@@
GPU Year
bfloat16
[FLOPs/s]
Memory Bandwidth
[Byte/s]
Arithmetic Intensity
[FLOPs/byte]
Communication
Bandwidth
[Byte/s]
@@
D.2 Prefill Stage: Time To First Token
@@
$\text{FLOPs}_{\text{algo}}$ can be calculated analytically given the FLOPs calculations provided in Appendix C.3, $\alpha_{\text{ eff}{{}}}$ and $\epsilon$ need to be fitted using the measured data. Exemplarily, we show the runtimes fitted for the measured TTFT in Figure 14 (Transfomer) and Figure 15 (xLSTM) for different model sizes. We fit $\alpha_{\text{ eff}{{}}}$ and $\epsilon$ per model configuration on TTFTs obtained under various combinations of batch sizes and prefill lengths. Our fits show excellent agreement between the predictions from our quantitative runtime model and the measured data. In Figure 16 we further show the quotient of the fitted $\alpha_{\text{ eff}{{}}}$ and the hardware parameter $\alpha_{\text{ acc}{{}}}$ for all model sizes. If $\alpha_{\text{ eff}{{}}}/\alpha_{\text{ acc}{{}}}=1$, the hardware would be perfectly utilized according to our model. We see that for both Transformers and xLSTM, the quotient increases, thus larger models utilize the hardware better. Furthermore, both models show relatively similar trends and magnitudes, indicating that the empirical measurement setup allowed for a fair comparison.
@@
D.3 Generation Stage: Step Time
@@
$\text{Bytes}_{\text{mem,algo}}$ can be calculated analytically given the MemOps calculations provided in Appendix C.4, $\beta_{\text{ eff}{{}}}$ and $\epsilon$ need to be fitted using the measured data. Furthermore, we found that the fit quality for Transformer further improved by fitting another constant that scales with the batch size. Exemplarily, we show the runtimes fitted for the measured step times in Figure 17 (Transformer) and Figure 18 (xLSTM) for different model sizes. Again, we find a very good agreement between the predictions from our quantitative runtime model and the measured data.
@@
Appendix E Model Configurations
In this section, we list the model hyperparameters and sizes of all training runs in Token/Param (Sec. E.1) and IsoFLOP (Sec. E.2) of the dataset for our scaling law study.
E.1 Model Sizes and Hyperparameters in Token/Param Configuration
@@
E.2 Model Sizes and Hyperparameters in IsoFLOP Configuration

### v2

@@ 7 Conclusion
Our study provides a systematic comparison of scaling behaviors between xLSTM and Transformer architectures. We show that xLSTMs are Pareto-dominant in training loss versus compute, maintain consistent power-law exponents in the overtraining regime, and scale more efficiently with context length due to their linear complexity. While our results suggest a universal relationship between performance and model size that applies to both compute-optimal Transformers and xLSTM models, we find that compute-optimal xLSTM models are larger than their Transformer counterparts and that the compute-optimal model size of xLSTMs is robust to variations in context length. During inference, xLSTM models achieve lower time to first tokens and generation step times than Transformer models of the same size. These results are well explained by our runtime model, which is grounded in theoretical FLOP and memory operation calculations and shows close agreement with the empirical data. Throughout all experiments, we find that the advantages of xLSTM grow with context length, both for training and inference characteristics, positioning xLSTM as a promising and scalable architecture for future language models.
Reproducibility Statement
We release the code to reproduce our experiments, the datasets of training runs as well as results for inference publicly upon acceptance to facilitate future research in this direction. The datasets of training runs have been obtained using the publicly available xLSTM 7B training repository (https://github.com/NX-AI/xlstm-jax) using the model configurations stated in Appendix D. Inference results have been obtained using the publicly available benchmarking pipeline for efficient xLSTM kernels (https://github.com/NX-AI/mlstm_kernels), more specifically, the model benchmarks, not those for individual kernels.
@@ A.1 Details on the Experimental Setup
In Appendix D we provide a list of model architecture configurations for all Transformer and xLSTM models used in our scaling law study in Token/Param (App. D.1) and IsoFLOP (App. D.2) training setups.
@@ A.4 Additional Results: IsoFLOP Approach
In the main paper (Sec. 3.4, Fig. 4), we presented results for the compute-optimal model size. In Fig. 9 we present results w.r.t. the number of training tokens. We observe that compute-optimal xLSTMs and Transformers are trained on a similar number of tokens.
@@ A.5 Additional Results: IsoFLOP Approach for Different Context Lengths
Appendix B Accounting: Parameters, Cache Sizes, FLOPs, Memory Operations
In this section, we count number of parameters (App. B.1), memory state or KV cache size B.2, FLOPs (App. B.3), and memory operations (App. B.4) for mLSTM models based on the architecture of xLSTM 7B (Beck et al., 2025b) and Transformer models with Self-Attention based on the Llama 3 architecture (Grattafiori et al., 2024).
@@ Appendix B Accounting: Parameters, Cache Sizes, FLOPs, Memory Operations
B.1 Parameter Counts
We count the number of parameters of mLSTM models (B.1.1) and Transformer models (B.1.2). We include embedding and normalization layer parameters in our parameter counts.
B.1.1 mLSTM Params
@@ B.1.2 Transformer Params
B.2 Memory State and KV-Cache Size
@@
B.3 FLOP Counts
In this section, we count the FLOPs for the mLSTM and the Transformer model architecture. For each model architecture we count the sequence length dependent FLOPs for the sequence mix layer first, i.e. the mLSTM cell (B.3.1) and the Self-Attention layer (B.3.3), and then combine them with the FLOPs of the other layers in the model architecture to obtain the total FLOPs for the mLSTM (B.3.2) and the Transformer model (B.3.4).
We do not drop subleading terms and set also count FLOPs for all operations equally, i.e. $F_{\text{OP}}=1$. We also count the FLOPs for the normalization layers with $F_{\text{norm}}=3$ (we assume the factor of 3 because we have mean, variance and division operations). The skip connection FLOPs are counted with $F_{\text{skip}}=1$, or if neglected with $F_{\text{skip}}=0$. Following our training configuration, we use the chunkwise-parallel formulation with chunk size $L=64$ and $F_{\text{causal}}=0.5$ for the FLOP counts and scaling laws in the main text.
B.3.1 mLSTM Cell FLOPs
@@ B.3 FLOP Counts
B.3.2 mLSTM Model FLOPs
The number of FLOPs for the backbone is identical for training, prefill and generation as the operations (embeddings, linear layers and layernorms) do not depend on the sequence length. Therefore, we count the FLOPs per token for the mLSTM backbone. To obtain the total FLOPs for the specific setting we have to use the respective expression for the mLSTM cell FLOPs from Appendix B.3.1.
@@
B.3.3 Self-Attention FLOPs
@@
B.3.4 Transformer Model FLOPs
Similar to the mLSTM backbone in Appendix B.3.2, the number of FLOPs for the Transformer backbone is identical for training, prefill and generation as the operations (embeddings, linear layers and layernorms) do not depend on the sequence length. Therefore, we count the FLOPs per token for the Transformer backbone. To obtain the total FLOPs for the specific setting we have to use the respective expression for the Self-Attention layer FLOPs from Appendix B.3.3.
@@ B.3.4 Transformer Model FLOPs
B.4 Memory Operation Counts
In this section, we count the memory operations for the mLSTM and the Transformer model architecture. We follow the same outline as for the FLOP counts in Appendix B.3 and first count the memory operations for the mLSTM cell (B.4.1) and the Self-Attention layer (B.4.3), and then combine them with the memory operations of the other layers in the model backbone to obtain the total memory operations for the mLSTM (B.4.2) and the Transformer model (B.4.4). We model weight MemOps as a one-time streaming cost (perfect on-chip reuse), i.e., independent of the number of token in the batch $BT$. This is reasonable with persistent/fused kernels and per-rank weight matrices that fit in on-chip cache. Depending on the exact experimental configuration, this assumption might not hold as we observe when modeling the step time through MemOps in Section C.3.
@@ B.4 Memory Operation Counts
B.4.1 mLSTM Cell MemOps
Similar to the FLOP counts in Appendix B.3.1, we count the memory operations for the mLSTM cell for both the chunkwise-parallel and the recurrent formulation.
@@
B.4.2 mLSTM Model MemOps
@@
B.4.3 Self-Attention MemOps
Similar to the FLOP counts in Appendix B.3.3, we count the memory operations for a single Self-Attention head during training or prefill and generation.
@@ B.4.3 Self-Attention MemOps
Similar to the FLOP counts in Appendix B.3.3, also the memory operation counts for the Self-Attention layer during generation depend on the number of previous tokens $T=T_{\text{p}}+t_{\text{g}}$, where $T_{\text{p}}$ is the number of prefill tokens and $t_{\text{g}}$ is the number of generated tokens so far.
@@
B.4.4 Transformer Model MemOps
Similar to the mLSTM backbone in Appendix B.4.2, the number of memory operations for the Transformer backbone (excluding the Self-Attention layer) consist of the input and output activations as well as the parameters. The memory operations for input and output activations depend on the number of tokens $BT$ in the batch, whereas the parameters are independent of the number of tokens.
@@ B.4.4 Transformer Model MemOps
Appendix C Modeling Inference Characteristics
In this section, we create a model of the theoretical runtimes of operations in the xLSTM and Transformer model architectures to model their inference characteristics (TTFT and step time). This theoretical model is based on the FLOP and the memory operation counts in Appendix B.
@@ Appendix C Modeling Inference Characteristics
C.1 Background: Theoretical Runtime
@@
For accelerator speed $\alpha_{\text{ acc}{{}}}$, accelerator memory bandwidth $\beta_{\text{ acc}{{}}}$, and accelerator communication bandwidth $\gamma_{\text{ Bytes}{{}}}$, we use the hardware specifications of NVIDIA V100333https://www.nvidia.com/en-au/data-center/v100/, A100444https://www.nvidia.com/en-us/data-center/a100/, H100555https://www.nvidia.com/en-au/data-center/h100/ and B200666https://resources.nvidia.com/en-us-blackwell-architecture/datasheet GPUs, which we summarize in Tab. 18.
@@
GPU Year bfloat16 [FLOPs/s] Memory Bandwidth [Byte/s] Arithmetic Intensity [FLOPs/byte] Communication Bandwidth [Byte/s]
@@
C.2 Prefill Stage: Time To First Token
@@
$\text{FLOPs}_{\text{algo}}$ can be calculated analytically given the FLOPs calculations provided in Appendix B.3, $\alpha_{\text{ eff}{{}}}$ and $\epsilon$ need to be fitted using the measured data. Exemplarily, we show the runtimes fitted for the measured TTFT in Figure 14 (Transfomer) and Figure 15 (xLSTM) for different model sizes. We fit $\alpha_{\text{ eff}{{}}}$ and $\epsilon$ per model configuration on TTFTs obtained under various combinations of batch sizes and prefill lengths. Our fits show excellent agreement between the predictions from our quantitative runtime model and the measured data. In Figure 16 we further show the quotient of the fitted $\alpha_{\text{ eff}{{}}}$ and the hardware parameter $\alpha_{\text{ acc}{{}}}$ for all model sizes. If $\alpha_{\text{ eff}{{}}}/\alpha_{\text{ acc}{{}}}=1$, the hardware would be perfectly utilized according to our model. We see that for both Transformers and xLSTM, the quotient increases, thus larger models utilize the hardware better. Furthermore, both models show relatively similar trends and magnitudes, indicating that the empirical measurement setup allowed for a fair comparison.
@@
C.3 Generation Stage: Step Time
@@
$\text{Bytes}_{\text{mem,algo}}$ can be calculated analytically given the MemOps calculations provided in Appendix B.4, $\beta_{\text{ eff}{{}}}$ and $\epsilon$ need to be fitted using the measured data. Furthermore, we found that the fit quality for Transformer further improved by fitting another constant that scales with the batch size. Exemplarily, we show the runtimes fitted for the measured step times in Figure 17 (Transformer) and Figure 18 (xLSTM) for different model sizes. Again, we find a very good agreement between the predictions from our quantitative runtime model and the measured data.
@@
Appendix D Model Configurations
In this section, we list the model hyperparameters and sizes of all training runs in Token/Param (Sec. D.1) and IsoFLOP (Sec. D.2) of the dataset for our scaling law study.
D.1 Model Sizes and Hyperparameters in Token/Param Configuration
@@
D.2 Model Sizes and Hyperparameters in IsoFLOP Configuration
@@
Appendix E Compute Optimal Parameter, Token and FLOP Count Estimates
In this section, we determine compute-optimal training setups for various model sizes based on the scaling laws derived from our IsoFLOP approach in sections 3.4 and 3.5. In Section E.1, we show the configurations for our power laws obtained for a context length of 8192 (see Figures 4 and 9), while in Section E.2 we present the compute optimal configurations obtained from our power law fits for varying context lengths (see Figures 5 and 10). The power law fits for Section E.2 contain fewer IsoFLOP profiles than the fits for Section E.1.
We construct these tables by first choosing a range of model sizes, then identifying the optimal compute budget associated with each size (for example, from Figures 4 or 5), and finally inferring the corresponding optimal number of training tokens, such as from Figures 9 or 10.
Across all tables in sections E.1 and E.2, we observe that Transformer models have a higher compute-optimal token-to-parameter ratio than xLSTM models.
Moreover, in contrast to the Chinchilla scaling laws, which find that the optimal token-to-parameter ratio is constant at around 22 across model sizes (Hoffmann et al., 2022, Table 3), our compute optimal token-to-parameter ratio decreases for larger models. This difference arises primarily from the distinct exponents in the scaling laws (Ours: $a=0.575$, $b=0.424$ vs. (Hoffmann et al., 2022, Table 2): $a=0.49(0.462,0.534)$, $b=0.51(0.483,529)$). Porian et al. (2024) have investigated these discrepancies and found the root cause to be in the learning rate decay for the training runs in the IsoFLOP configurations (see also Appendix A.3). They found exponents comparable to those in our work and were able to reproduce the Chinchilla scaling law exponents by using a fixed learning rate across all IsoFLOP training runs.
E.1 Compute Optimal Configurations for Context Length 8192
Table 23: Estimated optimal training FLOPs, Tokens, and Token/Param Ratio for varying model sizes from IsoFLOP power-law fits for Transformer and xLSTM models trained with context length 8192. The table is obtained from Figures 4 and 9.
Transformer xLSTM
#FLOPs $A^{\prime}=0.0023$ $a=0.575$ #Tokens $B^{\prime}=58.5$ $b=0.424$ Token/ Param Ratio #FLOPs $A^{\prime}=0.012$ $a=0.547$ #Tokens $B^{\prime}=77.7$ $b=0.417$ Token/ Param Ratio
#Params
100M 3.24e18 4.17B 41.7 1.33e18 2.83B 28.3
400M 3.61e19 11.6B 29.0 1.68e19 8.15B 20.4
1B 1.78e20 22.8B 22.8 8.97e19 16.4B 16.4
2B 5.94e20 38.1B 19.0 3.18e20 27.8B 13.9
4B 1.98e21 63.5B 15.9 1.13e21 47.1B 11.8
8B 6.62e21 106B 13.2 4.01e21 79.9B 10.0
10B 9.76e21 125B 12.5 6.03e21 94.8B 9.5
14B 1.75e22 160B 11.4 1.11e22 122B 8.7
32B 7.38e22 295B 9.2 5.05e22 230B 7.2
67B 2.67e23 508B 7.6 1.95e23 404B 6.0
175B 1.42e24 1.03T 5.9 1.13e24 840B 4.8
E.2 Compute Optimal Configurations for Varying Context Lengths
Table 24: Estimated optimal training FLOPs, Tokens, and Token/Param Ratio across context lengths from IsoFLOP context-specific power-law fits for Transformer models. The table is obtained from Figures 5 and 10.
ctx length: 2048 8192 16384 #FLOPs $A^{\prime}=0.0069$ $a=0.553$ #Tokens $B^{\prime}=74$ $b=0.423$ Token/ Param Ratio #FLOPs $A^{\prime}=0.0021$ $a=0.577$ #Tokens $B^{\prime}=65$ $b=0.422$ Token/ Param Ratio #FLOPs $A^{\prime}=0.0025$ $a=0.569$ #Tokens $B^{\prime}=34.5$ $b=0.432$ Token/ Param Ratio #Params 100M 2.27e18 4.24B 42.4 3.26e18 4.19B 41.9 4.19e18 3.91B 39.1 400M 2.78e19 12.2B 30.5 3.59e19 11.5B 28.8 4.78e19 11.2B 28.0 1B 1.46e20 24.6B 24.6 1.76e20 22.5B 22.5 2.39e20 22.5B 22.5 2B 5.09e20 41.7B 20.8 5.84e20 37.4B 18.7 8.07e20 38.1B 19.0 4B 1.78e21 70.8B 17.7 1.94e21 62.1B 15.5 2.73e21 64.5B 16.1 8B 6.23e21 120B 15.0 6.44e21 103B 12.9 9.21e21 109B 13.6 10B 9.32e21 143B 14.3 9.48e21 121B 12.1 1.36e22 129B 12.9 14B 1.71e22 184B 13.1 1.7e22 155B 11.1 2.46e22 167B 11.9 32B 7.62e22 346B 10.8 7.11e22 284B 8.9 1.05e23 313B 9.8 67B 2.9e23 609B 9.1 2.56e23 487B 7.3 3.85e23 549B 8.2 175B 1.64e24 1.27T 7.3 1.35e24 982B 5.6 2.08e24 1.14T 6.5
Table 25: Estimated optimal training FLOPs, Tokens, and Token/Param Ratio across context lengths from IsoFLOP context-specific power-law fits for xLSTM models. The table is obtained from Figures 5 and 10.
ctx length: 2048 8192 16384 #FLOPs $A^{\prime}=0.0086$ $a=0.555$ #Tokens $B^{\prime}=141$ $b=0.403$ Token/ Param Ratio #FLOPs $A^{\prime}=0.0161$ $a=0.541$ #Tokens $B^{\prime}=46.8$ $b=0.429$ Token/ Param Ratio #FLOPs $A^{\prime}=0.005$ $a=0.566$ #Tokens $B^{\prime}=336$ $b=0.385$ Token/ Param Ratio #Params 100M 1.32e18 2.83B 28.3 1.3e18 2.74B 27.4 1.58e18 3.44B 34.4 400M 1.6e19 7.73B 19.3 1.69e19 8.22B 20.6 1.83e19 8.85B 22.1 1B 8.32e19 15B 15.0 9.21e19 17B 17.0 9.23e19 16.5B 16.5 2B 2.9e20 24.9B 12.4 3.32e20 29.5B 14.8 3.14e20 26.5B 13.2 4B 1.01e21 41.1B 10.3 1.2e21 51B 12.8 1.07e21 42.4B 10.6 8B 3.51e21 68B 8.5 4.31e21 88.4B 11.0 3.64e21 68B 8.5 10B 5.25e21 79.9B 8.0 6.52e21 106B 10.6 5.39e21 79.1B 7.9 14B 9.62e21 102B 7.3 1.21e22 138B 9.9 9.77e21 99.5B 7.1 32B 4.26e22 186B 5.8 5.6e22 266B 8.3 4.21e22 175B 5.5 67B 1.61e23 318B 4.7 2.2e23 477B 7.1 1.55e23 289B 4.3 175B 9.07e23 638B 3.6 1.3e24 1.02T 5.8 8.47e23  [^5] B 3.2